OpenStudy (anonymous):

PLEASE HELP!!! (Algebra) y = 1/4x + 7 y = 1/2x + 5 FIND THE Y COORDINATE I have tried to solve this problem and this is what I have done: y = 1/4x + 7 and y = 1/2x + 5 Therefore… 1/4x + 7 = 1/2x +5 1/4x + 2 = 1/2x 2/8x + 2 = 4/8x 2 = 2/8x 1 = 1/4x 4 = x Now… ¼(4) + 7 = ½(4) + 5 1 + 7 = 2 + 5 8 DOES NOT EQUAL 7 What’s up? What did I do wrong? What is a better way to solve this equation?

OpenStudy (johnweldon1993):

Why did you need to find the least common denominator as 8? It'd be 4....dont forget 4 DOES go into 4 and so does 2 so here's what you had with corrections 1/4x + 7 = 1/2x +5 1/4x + 2 = 1/2x 1/4x + 2 = 2/4x 2 = 1/4x 2 = 1/4x 8 = x Now… ¼(8) + 7 = ½(8) + 5 2 + 7 = 4 + 5 9 DOES EQUAL 9 Hope that helped :)

OpenStudy (anonymous):

silly me, forgot that 4 :) Thank you so much!

OpenStudy (johnweldon1993):

haha yeah no problem...little mistakes happen to everyone :) glad I could help